Solve the following set of recurrence relations and initial conditions an = 4an−1 + n 2 − n , n ≥ 1; a0 = 1?


1
Expert's answer
2022-02-24T06:14:45-0500

Let us solve the following recurrence relation and initial condition

"a_n = 4a_{n−1} + n^2 − n ,\ n ≥ 1;\ a_0 = 1"

The characteristic equation "k-4=0" has the root "k=4," and hence the general solution of the reccurence relation is of the form


"a_n=A\cdot4^n+b_n," where the particular solution "b_n=pn^2+qn+r."


It follows that

"pn^2+qn+r=4(p(n-1)^2+q(n-1)+r)+n^2-n\\=4pn^2-8pn+4p+4qn-4q+4r+n^2-n."


Therefore,

"(3p+1)n^2+(-8p+3q-1)n+(4p-4q+3r)=0."

We conclude that

"3p+1=0,\ -8p+3q-1=0,\ 4p-4q+3r=0."

Then

"p=-\frac{1}3,\\ q=\frac{1}3(1+8p)=\frac{1}3(1-\frac{8}3)=-\frac{5}{9},
\\ r=\frac{4}3(q-p)=\frac{4}3(-\frac{5}9+\frac{1}3)=-\frac{8}{27}."


We conclude that the general solution of relation is the following:


"a_n=A\cdot4^n-\frac{1}3n^2-\frac{5}9n-\frac{8}{27}."


Since "a_0 = 1," we get "1=a_0=A-\frac{8}{27}." Then "A=1+\frac{8}{27}=\frac{35}{27}."


Consequently, the solution of the recurrence relation with initial condition is


"a_n=\frac{35}{27}\cdot4^n-\frac{1}3n^2-\frac{5}9n-\frac{8}{27}."